A solution contains thymol blue indicator. for eac
Answer #1

Thymol Blue gives following colours at the different pH

Thymol blue (pH indicator )                                     

                            below pH 1.2                    above pH 2.8                    below pH 8.0                        above pH 9.6

Colour             Brownish Red     \rightleftharpoons              Yellow                                  Yellow                                 Blue

a) \r\left [ H^{+} \right ]\, >\, \left [ OH^{-} \right ] \left [ H^{+} \right ]\, >\, \left [ OH^{-} \right ]             means the solution is acidic, so colour of the solution will be red/brownish red

b) \left [ H^{+} \right ]\, < \, \left [ OH^{-} \right ]                means the solution is basic, so the colour of solution will be blue

c) \left [ H^{+} \right ]\, = \, \left [ OH^{-} \right ]                means the solution is at neutral pH, (pH = 7). so the solution will give yellow colour

                                                      (pH 2.8 to 8.0 colour shown is yellow)